7, Earth has an orbital period of 365 days and its mean distance from the Su planet Pluto's mean distance from the Sun is 5.896x109 km. Using Kepler's the orbital period in Earth days.





8. The planet Venus orbits the Sun with a mean orbital radius of 1.076x101



is 1.99x1030 kg. Using Newton's version of Kepler's third law, calculate the orbital period of Venus.





1
Expert's answer
2021-12-02T10:04:11-0500

7) The orbital period of Pluto:


"T_P=T_E\sqrt{a_P^3/a_E^3}=90309."

8) The orbital period of Venus:


"T_V=2\pi\sqrt{\frac{a^3}{GM_\text{sun}}}=223\text{ days}."